# On Hardness and Approximation of Broadcasting in Sparse Graphs

### v1

@@ 3.1 Hardness of Telephone Broadcasting in $k$-cycle graphs
Given a multiset $C=\{c_{1},c_{2},\ldots,c_{m}\}$ of positive integers such that $\sum_{i=1}^{m}c_{i}=m(2m+1),$ does there exist a permutation $\alpha$ of $2,4,6,\ldots,2m$} and a permutation $\beta$ of $1,3,5,\ldots,2m-1$} such that $\alpha(i)+\beta(i)=c_{i}\quad\text{for all }1\leq i\leq m\,?$
@@
Consider the instance of Even-Odd RN3DM with $C=\{7,7,9,13\}$, where $m=4$. Note that $\sum_{i=1}^{4}c_{i}=36=m(2m+1)$. Now take the permutations $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$. One can check for every $i\in[4]$, $\alpha(i)+\beta(i)=c_{i}$. For instance $\alpha(1)+\beta(1)=4+3=7=c_{1}$, and $\alpha(2)+\beta(2)=6+1=7=c_{2}$. The same holds for $i=3,4$. Thus, $\alpha$ and $\beta$ certify that $C$ is a yes-instance of Even-Odd RN3DM.
@@
Figure 2: An instance $C=\{7,7,9,13\}$ of Even-Odd RN3DM reduces to the above instance of telephone broadcasting in $k$-cylce graphs (with $s$ as the originator and $br=8$). The broadcast tree associated with with solution $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$ is highlighted.
@@
Proof Sketch: We use a reduction from RN3DM, which is strongly NP-Hard [22]. Given an instance $W=\{w_{1},\ldots,w_{m}\}$ of RN3DM with target sum $e$, we construct an instance $C=\{c_{1},\ldots,c_{m}\}$ of Even-Odd RN3DM by setting $c_{i}=2e-2w_{i}-1\quad\text{for each }i\in[m].$ It is straightforward to verify that $W$ is a Yes-instance of RN3DM if and only if $C$ is a Yes-instance of Even-Odd RN3DM. The transformation uses unary encodings, so the reduction is strong. $\Box$
@@
(Reverse direction). Suppose $C$ is a yes-instance of Even-Odd RN3DM. Then there exist permutations $\alpha$ of $(2,4,\ldots,2m)$ and $\beta$ of $(1,3,\ldots,2m-1)$ such that for each $i\in[m]$, $\alpha(i)+\beta(i)=c_{i}.$ Define $\lambda(i)=\alpha(i)/2$ and $\mu(i)=(\beta(i)+1)/2$. Both $\lambda$ and $\mu$ are permutations of $[m]$. For each $i$,
@@
Thus $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
@@
Broadcasting in $k$-cycle graphs is NP-complete.
@@
Proof. Telephone Broadcasting is in NP in general graphs (and hence also for $k$-cycle graphs). We prove NP-hardness by a reduction from the Even-Odd RN3DM problem, which is strongly NP-Hard by Lemma 3.3.
@@
Let $C=\{c_{1},c_{2},\ldots,c_{m}\}$ be an instance of Even-Odd RN3DM. We construct an instance $(G,s,t)$ of Telephone Broadcasting in $k$-cycle graphs as follows. Graph $G$ consists of $m$ cycles, containing $c_{1},c_{2},\ldots,c_{m}$, extra vertices in addition to a single vertex $s$ that is shared between them. Now $s$ serves as the originator of the broadcast. Thus, the number of vertices in $G$ is $n=\sum_{i=1}^{m}c_{i}+1=m(2m+1)+1,$ and the degree of $s$ is $2m$. Since RN3DM is strongly NP-hard by Lemma 3.3, we may assume the values $c_{i}$ are encoded in unary, so the reduction is polynomial time. The decision problem asks whether broadcasting completes within $t=2m$ rounds. See Figure 2.
@@
(Forward direction). Suppose the answer to instance $C$ of Even-Odd RN3DM is yes. Then there exist permutations $\alpha$ of $(2,4,\ldots,2m)$ and $\beta$ of $(1,3,\ldots,2m-1)$ such that $\alpha(i)+\beta(i)=c_{i}$ for all $i\in[m]$. Construct a broadcast scheme as follows: the source $s$ informs its two neighbors in each cycle $c_{i}$ at times $2m-\alpha(i)+1$ and $2m-\beta(i)+1$, respectively. Given that $\alpha$ and $\beta$ are permutations of space $[2m]$, the source $s$ informs exactly one vertex at each given round $i\in[2m]$. Each newly informed vertex continues by informing its uninformed neighbor in the subsequent round. Hence, by round $br=2m$, exactly $\alpha(i)+\beta(i)=c_{i}$ vertices of the cycle (excluding $s$) are informed, and therefore the entire graph is informed. Thus, $(G,s,br)$ is a yes-instance of broadcasting.
@@
(Reverse direction). Now suppose $(G,s,br)$ is a yes-instance of broadcasting with $br=2m$. So, there is a broadcast scheme that completes within $22m$ rounds. Without loss of generality, assume the scheme is non-lazy, i.e., every informed vertex calls an uninformed neighbor whenever possible. Let $t_{i}$ denote the number of vertices newly informed in round $i$. At round 0, only $s$ is informed: $t_{0}=1$. At each round $i$, only $s$ and the vertices informed at time i1i-1 can inform new vertices (since other vertices of degree 2 have no additional uninformed neighbors left). Hence, $t_{i+1}\leq t_{i}+1$. It follows that the number of informed vertices by round $i$ is at most $1+1+2+\cdots+i=1+\frac{i(i+1)}{2}$. For $i=2m$, this reaches exactly $1+m(2m+1)=n$, so the broadcast must finish in exactly $2m$ rounds (and not earlier). This forces the source $s$ to make a call in every round, and since $\deg(s)=2m$, $s$ informs exactly one neighbor per round. Consequently, each cycle $c_{i}$ must be partitioned into two paths, informed from $s$ in opposite directions, of lengths $\alpha(i)$ and $\beta(i)$ with $\alpha(i)+\beta(i)=c_{i}$. Because $c_{i}$ is odd (otherwise the Even-Odd RN3DM instance would be trivial), we conclude that one of $\alpha(i),\beta(i)$ is odd and the other is even. This provides the exact decomposition required for a yes-instance of Even-Odd RN3DM. $\Box$
3.2 Hardness of Telephone Broadcasting in $k$-Path Graphs
@@
Next, we use a reduction from RN3DM to establish NP-completeness of broadcasting in $k$-path graphs.
@@
Broadcasting in $k$-path graphs is NP-complete.
Proof. Telephone Broadcasting belongs to NP for general graphs and hence also for $k$-path graphs. We prove NP-hardness by a reduction from the Restricted Numerical 3-Dimensional Matching (RN3DM) problem (Definition 3.1), which is strongly NP-Hard [22].
Let $W=\{w_{1},w_{2},\ldots,w_{m}\}$ be an instance of RN3DM. By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$. We construct an instance $(G,s,br)$ of Telephone Broadcasting in $k$-path graphs as follows. The graph $G$ consists of $m$ internally disjoint paths with shared endpoints $s$ and $t$, where the $i$-th path contains $\ell_{i}=e-w_{i}$ internal vertices in addition to $s$ and $t$. Furthermore, we add an edge connecting $s$ and $t$. The vertex $s$ serves as the source of the broadcast. The total number of vertices in $G$ is $n=2+\sum_{i=1}^{m}(e-w_{i})=2+me-\sum_{i=1}^{m}w_{i}=2+m(m+1).$ Both $s$ and $t$ have degree $2m+1$. Since RN3DM is strongly NP-hard, we may assume that the values $w_{i}$ are encoded in unary, ensuring that the reduction is polynomial time. The decision problem asks whether broadcasting completes within $br=m+1$ rounds (see Figure 3).
@@
Figure 3: An instance $W=\{1,3,4,4\}$ of RN3DM reduces to the above instance of telephone broadcasting in $k$-cylce graphs (with $s$ as the originator and $br=5$). The broadcast tree associated with with solution $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$ is highlighted.
4 Polynomial Time Approximation Schemes
4.1 PTAS for $k$-Cycle Graphs
In this section, we present a polynomial-time approximation scheme (PTAS) for Telephone Broadcasting in $k$-cycle graphs. Recall that a $k$-cycle graph is obtained by joining $k$ cycles at a common central vertex $r$. We assume that the number of cycles kk—and hence the optimal broadcast time—is asymptotically large; for constant $k$, the problem can be solved by a straightforward exhaustive search (see, e.g., [4]).
In what follows, we design a PTAS for the case where the broadcast originates at $s=r$. This restriction is without loss of generality: any $c$-approximation for the case that the center is the originator (when $s=r$) can be extended to a $c$-approximation in the general case ($s\neq r$). Suppose the originator $s$ is a non-center vertex on some cycle $C_{m}$, at distance $d\geq 1$ from $r$. The message can first be forwarded along the shortest path from $s$ to $r$. Once $r$ is informed, it first sends the message to its other neighbor on $C_{m}$ and then proceeds with the $c$-approximation algorithm on the rest of the graph. If the overall broadcast time is determined by a vertex in $C_{m}$, then the scheme is optimal. Otherwise, the process completes within $d+1+c\cdot\boldsymbol{br}^{-}$ rounds, where $\boldsymbol{br}^{-}$ is the optimal broadcast time excluding $C_{m}$. Since ${\boldsymbol{br}^{*}}\geq d+\boldsymbol{br}^{-}$, this yields the same approximation factor $c$.
We introduce the following problem, which is the key part of our PTAS algorithms.
@@ 4.1 PTAS for $k$-Cycle Graphs
The input to the Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers. The objective is to determine the minimum integer $m$ such that the set $C=[m]$ can be partitioned into $n$ subsets $C_{0},C_{1},\ldots,C_{n}$, such that for each $i\in[n]$, we have $|C_{i}|\leq 2$ and $\sum_{x\in C_{i}}x\,\geq\,\ell_{i}$. In this context, we say that $\ell_{i}$ is covered by the elements of $C_{i}$, and we refer to $C$ as the covering set of $S$.
@@
We begin by establishing the equivalence between Telephone Broadcasting in $k$-cycle graphs and the Prefix Covering problem.
@@
Let $G$ be a $k$-cycle graph with central vertex $s$, where the $i$-th cycle contains $c_{i}$ additional vertices besides $s$, and let $S=\{c_{1},\ldots,c_{k}\}$. Then Telephone Broadcasting for instance $(G,s)$ becomes equivalent to Prefix Covering for $S$.
@@
Therefore, Telephone Broadcasting in $k$-cycle graphs is equivalent to the Prefix Covering problem. $\Box$
4.1.1 PTAS for Prefix Covering
In this section, we present a PTAS for the Prefix Covering problem, that is, an algorithm with approximation factor $(1+\epsilon)$ for a given $\epsilon>1$. By Lemma 4.2, this would give a PTAS for Telephone Broadcasting in $k$-cycle graphs. As a high-level view of the algorithm, it addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity. The main idea is to transform the original ground multiset and the covering set into a simplified version with constant support (distinct elements), making it feasible to apply an exhaustive search which maintains provable approximation guarantees.
The algorithm operates as follows. Given an instance $S$ of Prefix Covering, we first construct a rounded set $\mathfrak{R}_{p}(S)$, where each element of $S$ is rounded up to a larger value in $S$ so that the support of $\mathfrak{R}_{p}(S)$ has size $p$. Here $p$ is a constant integer chosen as a function of $\epsilon$. Since rounding increases item values, covering $\mathfrak{R}_{p}(S)$ may require a larger covering set. Nevertheless, we show that the optimal covering of $\mathfrak{R}_{p}(S)$ is within a factor $(1+2/p)$ of the optimal covering of $S$ (Proposition 4.4). Thus, it suffices to focus on covering $\mathfrak{R}_{p}(S)$ rather than $S$. For that, we use a binary search over candidate values $m$, where $m$ specifies the size of the covering set $C=[m]$. For each $m$, we define a rounded version $\mathfrak{R}_{p}(C)$, obtained by rounding up the elements of $C$ so that its support also has size $p$. We prove that if $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}(C)$, then it can also be covered using $C^{\prime}=[m^{\prime}]$ with $m^{\prime}\leq(1+1/p)m$ (4.1). Conversely, if $\mathfrak{R}_{p}(S)$ cannot be covered by $\mathfrak{R}_{p}(C)$, then it cannot be covered by $C(=[m])$. Hence, verifying whether $\mathfrak{R}_{p}(C)$ successfully covers $\mathfrak{R}_{p}(S)$ is sufficient for approximating the original instance. This verification can be performed by exhaustive search over $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$. Since both sets have constant support (bounded by $p$), this step can be carried out in time independent of $n$ (Lemma 4.5). In what follows, we provide the details of the algorithm.

### v2

@@ 3.1 Hardness of Telephone Broadcasting in Cycle-star Graphs
Given a multiset $C=\{c_{1},c_{2},\ldots,c_{m}\}$ of positive integers such that $\sum_{i=1}^{m}c_{i}=m(2m+1),$ does there exist a permutation $\alpha$ of $\{2,4,6,\ldots,2m\}$ and a permutation $\beta$ of $\{1,3,5,\ldots,2m-1\}$ such that $\alpha(i)+\beta(i)=c_{i}\quad\text{for all }1\leq i\leq m\,?$
@@
Consider the instance of Even-Odd RN3DM with $C=\{7,7,9,13\}$, where $m=4$. Note that $\sum_{i=1}^{4}c_{i}=36=m(2m+1)$. Now take the permutations $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$. One can check for every $i\in[4]$, $\alpha(i)+\beta(i)=c_{i}$. For instance $\alpha(1)+\beta(1)=4+3=c_{1}$, and $\alpha(2)+\beta(2)=6+1=c_{2}$. The same holds for $i=3,4$. Thus, $\alpha$ and $\beta$ certify that $C$ is a yes-instance of Even-Odd RN3DM.
@@
Figure 2: An instance $C=\{7,7,9,13\}$ of Even-Odd RN3DM reduces to the above instance of Telephone Broadcasting in cycle-star graphs (with $s$ as the originator and $br=8$). The broadcast tree associated with solution $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$ is highlighted.
@@
Proof sketch: We use a reduction from RN3DM, which is strongly NP-hard [29]. Given an instance $W=\{w_{1},\ldots,w_{m}\}$ of RN3DM with target sum $e$, we construct an instance $C=\{c_{1},\ldots,c_{m}\}$ of Even-Odd RN3DM by setting $c_{i}=2e-2w_{i}-1\quad\text{for each }i\in[m].$ It is straightforward to verify that $W$ is a Yes-instance of RN3DM if and only if $C$ is a Yes-instance of Even-Odd RN3DM. $\Box$
@@
(Reverse direction). Suppose $C$ is a yes-instance of Even-Odd RN3DM. Then there exist permutations $\alpha$ of $\{2,4,\ldots,2m\}$ and $\beta$ of $\{1,3,\ldots,2m-1\}$ such that for each $i\in[m]$, $\alpha(i)+\beta(i)=c_{i}.$ Define $\lambda(i)=\alpha(i)/2$ and $\mu(i)=(\beta(i)+1)/2$. Both $\lambda$ and $\mu$ are permutations of $[m]$. For each $i$,
@@
Thus, $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
@@
We are ready to prove the main result of this section, namely, the hardness of Telephone Broadcasting in cycle-star.
@@
Telephone Broadcasting in cycle-star graphs is NP-Complete.
@@
Proof. Telephone Broadcasting is in NP in general graphs (and hence also for cycle-star graphs). We prove NP-hardness by a reduction from the Even-Odd RN3DM problem, which is strongly NP-hard by Lemma 3.3.
@@
Let $C=\{c_{1},c_{2},\ldots,c_{m}\}$ be an instance of Even-Odd RN3DM. We construct an instance $(G,s,t)$ of Telephone Broadcasting in cycle-star graphs as follows. Graph $G$ consists of $m$ cycles, containing $c_{1},c_{2},\ldots,c_{m}$, extra vertices in addition to a single vertex $s$ that is shared between them. Now $s$ serves as the originator of the broadcast. Thus, the number of vertices in $G$ is $n=\sum_{i=1}^{m}c_{i}+1=m(2m+1)+1,$ and the degree of $s$ is $2m$. Since RN3DM is strongly NP-hard by Lemma 3.3, we may assume the values $c_{i}$ are encoded in unary, so the reduction is polynomial time. The decision problem asks whether broadcasting completes within $t=2m$ rounds. See Figure 2.
@@
(Forward direction). Suppose the answer to instance $C$ of Even-Odd RN3DM is yes. Then there exist permutations $\alpha$ of $\{2,4,\ldots,2m\}$ and $\beta$ of $\{1,3,\ldots,2m-1\}$ such that $\alpha(i)+\beta(i)=c_{i}$ for all $i\in[m]$. Construct a broadcast scheme as follows: the source $s$ informs its two neighbors in each cycle $c_{i}$ at times $2m-\alpha(i)+1$ and $2m-\beta(i)+1$, respectively. Given that $\alpha$ and $\beta$ are permutations of space $[2m]$, the source $s$ informs exactly one vertex at each given round $i\in[2m]$. Each newly informed vertex continues by informing its uninformed neighbor in the subsequent round. Hence, by round $br=2m$, exactly $\alpha(i)+\beta(i)=c_{i}$ vertices of the cycle (excluding $s$) are informed, and, therefore, the entire graph is informed. Thus, $(G,s,br)$ is a yes-instance of Telephone Broadcasting.
@@
(Reverse direction). Now suppose $(G,s,br)$ is a yes-instance of Telephone Broadcasting with $br=2m$. Therefore, there is a broadcast scheme that completes within $2m$ rounds. Without loss of generality, assume the scheme is non-lazy, i.e., every informed vertex calls an uninformed neighbor whenever possible. Let $t_{i}$ denote the number of vertices newly informed in round $i$. At round 0, only $s$ is informed: $t_{0}=1$. At each round $i$, only $s$ and the vertices informed at time i1i-1 can inform new vertices (since other vertices of degree 2 have no additional uninformed neighbors left). Hence, $t_{i+1}\leq t_{i}+1$. It follows that the number of informed vertices by round $i$ is at most $1+1+2+\cdots+i=1+\frac{i(i+1)}{2}$. For $i=2m$, this reaches exactly $1+m(2m+1)=n$, so the broadcast must finish in exactly $2m$ rounds (and not earlier). This forces the source $s$ to make a call in every round, and since $\deg(s)=2m$, $s$ informs exactly one neighbor per round. Consequently, each cycle $c_{i}$ must be partitioned into two paths, informed from $s$ in opposite directions, of lengths $\alpha(i)$ and $\beta(i)$ with $\alpha(i)+\beta(i)=c_{i}$. Because $c_{i}$ is odd (otherwise the Even-Odd RN3DM instance would be trivial), we conclude that one of $\alpha(i),\beta(i)$ is odd and the other is even. This provides the exact decomposition required for a yes-instance of Even-Odd RN3DM. $\Box$
3.2 Hardness of Telephone Broadcasting in Melon Graphs
@@
Next, we use a reduction from RN3DM to establish NP-Completeness of Telephone Broadcasting in melon graphs.
@@
Telephone Broadcasting in melon graphs is NP-Complete.
Proof sketch: We prove NP-hardness by a reduction from the Restricted Numerical 3-Dimensional Matching (RN3DM) problem (Definition 3.1), which is strongly NP-hard [29].
Let $W=\{w_{1},w_{2},\ldots,w_{m}\}$ be an instance of RN3DM. By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=m\cdot e$ for some integer $e$. We construct an instance $(G,s,br)$ of Telephone Broadcasting in melon graphs as follows. The graph $G$ consists of $m$ internally disjoint paths with shared endpoints $s$ and $t$, where the $i$-th path contains $\ell_{i}=e-w_{i}$ internal vertices in addition to $s$ and $t$. Furthermore, we add an edge connecting $s$ and $t$. The vertex $s$ serves as the source of the broadcast. The total number of vertices in $G$ is $n=2+\sum_{i=1}^{m}(e-w_{i})=2+me-\sum_{i=1}^{m}w_{i}=2+m(m+1).$ Both $s$ and $t$ have degree $2m+1$. Since RN3DM is strongly NP-hard, we may assume that the values $w_{i}$ are encoded in unary, ensuring that the reduction is polynomial time. The decision problem asks whether broadcasting completes within $br=m+1$ rounds (see Figure 3). The answer to this decision problem is yes if and only if the $W$ is a yes-instance of RN3DM. $\Box$
Proof. Telephone Broadcasting belongs to NP for general graphs and hence also for melon graphs. We prove NP-hardness by a reduction from the Restricted Numerical 3-Dimensional Matching (RN3DM) problem (Definition 3.1), which is strongly NP-hard [29].
Let $W=\{w_{1},w_{2},\ldots,w_{m}\}$ be an instance of RN3DM. By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=m\cdot e$ for some integer $e$. We construct an instance $(G,s,br)$ of Telephone Broadcasting in melon graphs as follows. The graph $G$ consists of $m$ internally disjoint paths with shared endpoints $s$ and $t$, where the $i$-th path contains $\ell_{i}=e-w_{i}$ internal vertices in addition to $s$ and $t$. Furthermore, we add an edge connecting $s$ and $t$. The vertex $s$ serves as the source of the broadcast. The total number of vertices in $G$ is $n=2+\sum_{i=1}^{m}(e-w_{i})=2+me-\sum_{i=1}^{m}w_{i}=2+m(m+1).$ Both $s$ and $t$ have degree $2m+1$. Since RN3DM is strongly NP-hard, we may assume that the values $w_{i}$ are encoded in unary, ensuring that the reduction is polynomial time. The decision problem asks whether broadcasting completes within $br=m+1$ rounds (see Figure 3).
@@
Figure 3: An instance $W=\{1,3,4,4\}$ of RN3DM reduces to the above instance of Telephone Broadcasting in cycle-star graphs (with $s$ as the originator and $br=5$). The broadcast tree associated with with solution $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$ is highlighted.
4 Efficient Polynomial Time Approximation Schemes
4.1 EPTAS for Cycle-star Graphs
In this section, we present an efficient polynomial-time approximation scheme (EPTAS) for Telephone Broadcasting in cycle-star graphs. Recall that a cycle-star graph is obtained by joining $k$ cycles at a common central vertex $r$. We assume that the number of cycles kk—and hence the optimal broadcast time—is asymptotically large; for constant $k$, the problem can be solved optimally by a straightforward exhaustive search (see, e.g., [5]).
In what follows, we design an EPTAS for the case where the broadcast originates at $s=r$. This restriction is without loss of generality: any $c$-approximation for the case that the center is the originator (when $s=r$) can be extended to a $c$-approximation in the general case ($s\neq r$). Suppose the originator $s$ is a non-center vertex on some cycle $Y$, at distance $d\geq 1$ from $r$. The message can first be forwarded along the shortest path from $s$ to $r$. Once $r$ is informed, it first sends the message to its other neighbor on $Y$ and then proceeds with the $c$-approximation algorithm on the rest of the graph. If the overall broadcast time is determined by a vertex in $Y$, then the scheme is optimal. Otherwise, the process completes within $d+1+c\cdot\boldsymbol{br}^{-}$ rounds, where $\boldsymbol{br}^{-}$ is the optimal broadcast time excluding $Y$. Since ${\boldsymbol{br}^{*}}\geq d+\boldsymbol{br}^{-}$, this yields the same approximation factor $c$.
From Telephone Broadcasting in Cycle-star Graphs to Prefix Covering
We introduce the following problem, which is the key part of our EPTAS algorithms.
@@ 4.1 EPTAS for Cycle-star Graphs
The input to the Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers. The objective is to determine the minimum integer $m$ such that the set $C=[m]$ can be partitioned into $n+1$ subsets $C_{0},C_{1},\ldots,C_{n}$, such that for each $i\in[n]$, we have $|C_{i}|\leq 2$ and $\sum_{x\in C_{i}}x\,\geq\,\ell_{i}$. In this context, we say that $\ell_{i}$ is covered by the elements of $C_{i}$, and we refer to $C$ as the covering set of $S$.
@@
We begin by establishing the equivalence between Telephone Broadcasting in cycle-star graphs and the Prefix Covering problem.
@@
Let $G$ be a cycle-star graph with central vertex $s$, where the $i$-th cycle contains $c_{i}$ additional vertices besides $s$, and let $S=\{c_{1},\ldots,c_{k}\}$. Then Telephone Broadcasting for instance $(G,s)$ becomes equivalent to Prefix Covering for $S$.
@@
Therefore, Telephone Broadcasting in cycle-star graphs is equivalent to the Prefix Covering problem. $\Box$
EPTAS for Prefix Covering
In this section, we present an EPTAS for the Prefix Covering problem, that is, an algorithm with approximation factor $(1+\epsilon)$ for a given $\epsilon>0$. By Lemma 4.2, this would give an EPTAS for Telephone Broadcasting in cycle-star graphs. As a high-level view of the algorithm, it addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity. The main idea is to transform the original ground multiset and the covering set into a simplified version with constant support (i.e., distinct elements), enabling formulation as a feasibility ILP with a constant number of variables. Solving this ILP yields to the desired EPTAS.
The algorithm operates as follows. Given an instance $S$ of Prefix Covering, we first construct a rounded set $\mathfrak{R}_{p}(S)$, where each element of $S$ is rounded up to a larger value in $S$ so that the support of $\mathfrak{R}_{p}(S)$ has size $p$. Here $p$ is a constant integer chosen as a function of $\epsilon$. Since rounding increases item values, covering $\mathfrak{R}_{p}(S)$ may require a larger covering set. Nevertheless, we show that the optimal covering of $\mathfrak{R}_{p}(S)$ is within a factor $(1+2/p)$ of the optimal covering of $S$ (Proposition 4.4). Thus, it suffices to focus on covering $\mathfrak{R}_{p}(S)$ rather than $S$. For that, we use a binary search over candidate values $m$, where $m$ specifies the size of the covering set $C=[m]$. For each $m$, we define a rounded version $\mathfrak{R}_{p}(C)$, obtained by rounding up the elements of $C$ so that its support also has size $p$. We prove that if $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}(C)$, then it can also be covered using $C^{\prime}=[m^{\prime}]$ with $m^{\prime}\leq(1+1/p)m$ (4.1). Conversely, if $\mathfrak{R}_{p}(S)$ cannot be covered by $\mathfrak{R}_{p}(C)$, then it cannot be covered by $C(=[m])$. Hence, verifying whether $\mathfrak{R}_{p}(C)$ successfully covers $\mathfrak{R}_{p}(S)$ is sufficient for approximating the original instance. This verification can be performed by solving an ILP over $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$. Since both sets have constant support (bounded by $p$), this step can be carried out in time independent of $n$ (Lemma 4.5). In what follows, we provide the details of the algorithm.